OpenStudy (anonymous):

f(x) = 10 x^2-9 find domain and range

OpenStudy (danjs):

are there any values for x that won't work?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

sorry i have no idea on where to start

OpenStudy (danjs):

nah, x=9 will return a value for f(x) THe domain is all the input values that work the range is the values f(x) takes

OpenStudy (danjs):

For x^2, any real value will work, right

OpenStudy (anonymous):

right

OpenStudy (danjs):

THe parabola opens in the vertical direction, it gets wider and wider forever, the values for x can be any real value That is the domain

OpenStudy (anonymous):

so its infinity?

OpenStudy (danjs):

-infinity to +infinity

OpenStudy (jdoe0001):

\(\Large (-\infty,+\infty)\)

OpenStudy (danjs):

THe domain of f is all values for x, such that, x is an element of the reals

OpenStudy (anonymous):

so the range would also be (infinity,-infinity)

OpenStudy (danjs):

hmm, look at the function. x^2 is the only variable, x^2 will always be posistive or zero

OpenStudy (jdoe0001):

he's referring to the domain ony thus far =) \(\large \{x|\quad x\in \mathbb{R}\}\)

OpenStudy (danjs):

how does that affect what values f(x) can take

OpenStudy (anonymous):

so x would be 9?

OpenStudy (danjs):

no x will be any real number, from the domain we just figured f(x) may be limited to some interval(s) sometimes, depending on the nature of the function

OpenStudy (danjs):

If x^2 is 0 or larger, how does that limit the values for f(x) ?

OpenStudy (danjs):

something squared will be positive, (-1)^2 = -1 * -1 = +1

OpenStudy (danjs):

To graph this parabola, you probably will figure the vertex point, and the + on the ax^2 term means it opens toward the +y direction

OpenStudy (anonymous):

right I got the graph

OpenStudy (danjs):

if x^2 is always 0 or larger, the smallest value for f(x) is when x = 0 , and f(x) = -9, and the thing opens towards positive y. the values for f(x), THE RANGE , is then limited to y values greater than or equal to -9

OpenStudy (anonymous):

so in interval notation would it look like (-9,infinity)

OpenStudy (danjs):

bra-ket for the -9, it can be that value

OpenStudy (danjs):

greater than or equal to -9

OpenStudy (danjs):

[-9 , inf)

OpenStudy (anonymous):

oh my gosh, okay thank you, im starting to get it now

OpenStudy (danjs):

welcome, takes practicing